OpenStudy (anonymous):

Using the definition of the derivative find the equation of the line tangent to the curve at x=3 if f(x)=2x^2-13x+5

OpenStudy (anonymous):

Take the derivative, what do you get?

OpenStudy (anonymous):

4x-13

OpenStudy (anonymous):

Precisely. Now evaluate it at x=3 and you will have the slope of your tangent line. To solve for the equation of the line you'll need: \[y-y_0=m(x-x_0)\] where m is the slope we just found and (x_0,y_0) is an arbitrary point on the line. In this case, use x=3 and evaluate f(3). From this you have your point (x_0,y_0). Use this in the above formula for a line with the slope,m, given by f'(3).

OpenStudy (anonymous):

Hi Fudger The general way that you would do this is to find the instantaneous rate of change at that point by using the derivative and also find the coordinates of x=3 by using the original equation. You can then substitute that rate of change and the coordinate into a linear equation (which I forgot the name of) in order to produce an equation with a line going through that coordinate, with the gradient that we found earlier. This happens to be the tangent. So, find the instantaneous rate of change: f(x) = 2x^2-13x+5 f'(x) = 4x-13 (derivative) f'(3) = 4*3-13 f'(3) = -1 Therefore the gradient at that point is -1. Now find the original coordinate: f(x) = 2x^2-13x+5 f(3) = 2*(3^2)-13*3+5 f(3) = 18-39+5 f(3) = -16 Therefore the coordinate of the point is (3, -16). Now use the linear formula to calculate the equation for the tangent: y-y1 = m(x-x1) (where m is the gradient) y+16 = -1(x-3) y+16 = -x+3 y = -x - 13 Therefore the answer to your question is y = -x-13. You can check this in wolfram alpha: http://goo.gl/5XDDrE And you can see that the line is a tangent.

OpenStudy (anonymous):

So, f(3) would be -16. So you need: \[(x_0,y_0)=(3,-16)\]

OpenStudy (anonymous):

thank you!